# Urban Region Representation Learning: A Flexible Approach

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Location based servicesccs: Information systems Data mining
@@ 1. Introduction
Urban region representation learning has become an increasingly popular topic in the community of urban computing (Zhang et al., 2017; Zheng et al., 2014; Yuan et al., 2024; Liu et al., 2024; Gupta et al., 2023), which aims to transform urban regions into vector representations, known as embeddings. These embeddings entail valuable insights on urban structures and properties, facilitating effective urban planning and management, such as designating functionalities for new development areas. They are also useful in various tasks related to daily life, such as crime count prediction (Wu et al., 2022; Sun et al., 2024; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024).
@@
Limitation 1. Existing methods heavily depend on mobility data and underutilize publicly accessible data. Human mobility data plays a critical role in learning effective region embeddings (Sun et al., 2024; Wu et al., 2022; Li et al., 2024; Zhang et al., 2020, 2023b). However, their limited availability, particularly in underdeveloped regions, together with privacy issues, prevent models using such data from a wider adoption. Recently, several studies (Li et al., 2023; Xi et al., 2022; Yan et al., 2024; Yong and Zhou, 2024) leverage features from publicly accessible data (e.g., POIs from OpenStreetMap or street view images from Google Maps) to enhance model applicability. However, these studies suffer from the effectiveness of the learning models. Their learned embeddings have reported lower accuracy for downstream tasks comparing with those learned by the mobility-based models.
@@
We summarize existing works for the issues above in Table 1. A detailed discussion on these works can be found in Appendix A.
@@
(2) It has a prompt learning process for its third stage, which enables it to flexibly utilize different types of features for different downstream tasks. We propose a novel prompt enhancer (PromptEnhancer) model to tailor region embeddings for downstream tasks by integrating textual and street-view imagery features. To capture task-relevant information, PromptEnhancer consists of a text-region alignment (T-RAlign) module and a street view-region alignment (SV-RAlign) module. T-RAlign incorporates task-specific geographic knowledge into region embeddings using dimension-wise similarity, while SV-RAlign extracts task-relevant visual features through adapted attention mechanisms. (Addressing Limitation 2)
@@ 2. Solution Overview
Model overview. Fig. 2 shows the overall structure of our model FlexiReg, which consists of three main learning stages. (1) FlexiReg takes a set of grid cells as the input. The grid cells come from a fine-grained partitioning over the spatial area of interest that we perform as part of data preparation (Section 3.1). FlexiReg learns the embeddings of grid cells across different features through a Multimodal Grid Cell Embedding Learning module (GridLearner, Section 3.2). (2) Then, the Adaptive Region Embedding Generation module (AdaRegionGen) aggregates the fine-grained cell embeddings to generate region embeddings for the input regions (Section 3.3). Starting from the fine-grained cell embeddings allows FlexiReg to flexibly adapt to different sets of regions which may come from different downstream tasks (or analytical tasks to explore different way to form regions). (3) Finally, the Prompt Enhancer module (PromptEnhancer) refines the generic region embeddings with extra features, guided by and tailored for given downstream tasks (Section 3.4).
@@ 3. Proposed Model
This section details the FlexiReg model. We include a notation table in Section B.
3.1. Data Preparation
3.1.1. Grid Cell Construction
3.1.2. Feature Preparation
We use six types of features for each cell, which are all publicly accessible, with full details in Appendix C.
POI features. For each cell, we count the number of POIs that belong to one of 15 POI categories from OpenStreetMap (OpenStreetMap, 2024) as the POI feature. We denote the POI feature of cell $c_{i}$ as $\mathbf{p}_{i}\in\mathbb{R}^{15}$.
Land use features. Similar to POI features, we count the numbers of zones that belong to 20 different land use types within a cell. The land use feature of cell $c_{i}$ is denoted as $\mathbf{l}_{i}\in\mathbb{R}^{20}$.
Street view imagery features. Street view images capture finer-grained urban patterns. We use $\mathbf{sv}_{i}=\{\mathbf{sv}_{i,1},\mathbf{sv}_{i,2},\cdots\}$ to denote the set of street view images captured within the area of $c_{i}$, where each image is also in the shape of $\mathbb{R}^{H\times W\times 3}$. Note that different cells may have different numbers of street view images.
3.2. Multimodal Grid Cell Embedding Learning
3.2.1. Grid-based Intra-view Feature Learning
(1) $a_{ij}^{g}$ $=\sigma\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{g}_{i}||\mathbf{W}\mathbf{z}^{g}_{j}||\mathbf{w}\mathbf{A}_{i,j}\right)\right),$
(2) $\alpha_{ij}^{g}$ $=\mathrm{Softmax}(a_{ij}^{g}),$
(4) $\mathbf{z}^{si}_{i}=\mathrm{MLP}(\mathrm{ResNet}(\mathbf{si}_{i})),$
3.2.2. Grid-based Inter-view Feature Learning
Next, we learn the correlation between different feature views for each grid cell by applying a one-layer self-attention (Vaswani et al., 2017) on $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$, which is obtained by stacking $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$ and $\mathbf{Z}^{si}$. We update $\mathbf{Z}$ by adaptive aggregation to be the output of self-attention. More details of this module can be found in Appendix D.
3.2.3. Grid-based Dual-Feature Attentive Fusion
We adopt the dual-feature attentive fusion module (DAFusion) from HAFusion (Sun et al., 2024) to further refine cell embeddings $\mathbf{Z}$. DAFusion first fuses the representations among different views into an adaptive view representation for the cells. Then, it fuses the representations among cells. We denote the output of this module as $\mathbf{E}\in\mathbb{R}^{m\times d}$. Due to space limit, we elaborate this module in Appendix E.
3.2.4. Module Training.
POI-oriented and land use-oriented objectives $\mathcal{L}^{p}$ and $\mathcal{L}^{l}$. We use graph reconstruction as the learning objective to reconstruct the POI adjacency matrix $\mathbf{A}^{p}$ and land use adjacency matrix $\mathbf{A}^{l}$ from their feature-oriented embeddings $\mathbf{E}^{p}$ and $\mathbf{E}^{l}$, respectively. We leave the details in Appendices F.1 and F.2.
Geographic neighbor-oriented objective $\mathcal{L}^{gn}$. Inspired by the First Law of Geography (Tobler, 1970), emphasizing that spatially close cells are likely to share similar functionality and embeddings, we use a triplet loss (Schroff et al., 2015) to learn geographic neighbor-oriented embeddings. We leave the details in Appendix F.3.
Satellite image-oriented objective $\mathcal{L}^{si}$. We model the satellite image-oriented objective as predicting the number of POIs in a cell using the corresponding satellite image, where the ground-truth POI count can be obtained from the POI feature. Such a training objective guides our model to focus on spatial characteristics when learning embeddings from satellite images. The details of this objective can be found in Appendix F.4.
3.3. Adaptive Region Embedding Generation
Here, $\mathbf{e}_{i}\in\mathbf{E}$ is the cell embedding of $c_{i}$.
3.4. Prompt Enhancer for Task Learning
3.4.1. Text-Region Encoding
An example of the generated textual description of a cell can be found in Appendix G.
We use the same cell-to-region embedding aggregation approach as described in Section 3.3 to obtain the region embeddings from textual features, i.e., following Equations 6 and  7. The textual embeddings of regions are denoted as $\mathbf{H}^{t}\in\mathbb{R}^{n\times d_{llm}}$.
Figure 3. Text-region alignment module
3.4.2. Street View-Region Encoding
Figure 4. Street view image embedding learning
(10) $\mathbf{u}_{i,j}=\mathrm{ResNet}(\mathbf{sv}_{i,j}).$
where $|\mathbf{sv}_{i}|$ is the number of images associated with $c_{i}$.
3.4.3. Model Training.
where $y_{i}$ is the ground truth for $r_{i}$, and $n$ is the number of regions.
4. Experiments
We run experiments to verify: (Q1) the embedding quality of our FlexiReg model as compared with the state-of-the-art (SOTA) models on four downstream tasks, (Q2) the applicability of FlexiReg across diverse geographic regions, (Q3) the adaptability of our cell embeddings to different region formations, (Q4) the impact of our model components and input features, (Q5) the impact of urban environments, the grid cell design, and key hyper-parameters.
4.1. Experimental Settings
Dataset. We use real data from five cities across the globe: New York City (NYC(New York Dataset, 2020), Chicago (CHI(Chicago Dataset, 2020), San Francisco (SF(San Francisco Dataset, 2020), Singapore City (SG(Singapore Dataset, 2019), and Lisbon (LX(Lisbon Region Boundary, 2024). We collect data on region division, POI, land use, satellite images, street view images, check-in, and population. Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities). More details about the collected data can be found in Appendix H.
Competitors. We compare with models from two categories. The first category uses a subset of the publicly accessible data: RegionDCL (Li et al., 2023), UrbanCLIP (Yan et al., 2024), and CityFM (Balsebre et al., 2024) (SOTA). The second category uses human mobility data, which has restricted availability: MVURE (Zhang et al., 2020), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023), ReCP (Li et al., 2024), and HAFusion (Sun et al., 2024) (SOTA). This latter category does not apply to SG and LX due to lack of data. Detail descriptions of these models are included in Appendix I.
Detailed implementation and hyperparameter settings of these models and our FlexiReg can be found in Appendix J.
We evaluate the representation learning models through the downstream models in mean absolute error (MAE), root mean square error (RMSE), and coefficient of determination ($R^{2}$).
4.2. Overall Results (Q1)
The overall model accuracy results are reported in Table 2. Here, we only report $R^{2}$ for conciseness, as the performance in MAE and RMSE resembles. The full results can be found in Appendix K.
We make the following observations.
Table 2. Overall Prediction Accuracy Results ($\uparrow$’ indicates that large values are preferred. The best results are in boldface, and the second-best results are underlined. Same for the tables below.)
New York City Crime Check-in Service Call Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 0.591 $\pm$ 0.016 0.627 $\pm$ 0.019 0.367 $\pm$ 0.027 0.545 $\pm$ 0.008
MGFN (Wu et al., 2022) 0.630 $\pm$ 0.020 0.690 $\pm$ 0.040 0.303 $\pm$ 0.069 0.509 $\pm$ 0.026
HREP (Zhou et al., 2023) 0.680 $\pm$ 0.014 0.703 $\pm$ 0.021 0.398 $\pm$ 0.021 0.571 $\pm$ 0.021
ReCP (Li et al., 2024) 0.459 $\pm$ 0.022 0.761 $\pm$ 0.029 0.356 $\pm$ 0.029 0.322 $\pm$ 0.047
HAFusion (Sun et al., 2024) 0.734 $\pm$ 0.015 0.844 $\pm$ 0.012 0.493 $\pm$ 0.014 0.616 $\pm$ 0.019
RegionDCL (Li et al., 2023) 0.251 $\pm$ 0.026 0.471 $\pm$ 0.023 0.103 $\pm$ 0.026 0.198 $\pm$ 0.019
UrbanCLIP (Yan et al., 2024) 0.267 $\pm$ 0.012 0.458 $\pm$ 0.005 0.232 $\pm$ 0.005 0.276 $\pm$ 0.002
CityFM (Balsebre et al., 2024) 0.315 $\pm$ 0.010 0.471 $\pm$ 0.011 0.117 $\pm$ 0.013 0.261 $\pm$ 0.002
FlexiReg 0.789 $\pm$ 0.009 0.876 $\pm$ 0.006 0.601 $\pm$ 0.021 0.715 $\pm$ 0.010
Improvement 7.5% 3.8% 21.9% 16.1%
Chicago Crime Check-in Service Call Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 0.461 $\pm$ 0.062 0.656 $\pm$ 0.029 0.441 $\pm$ 0.050 0.313 $\pm$ 0.043
MGFN (Wu et al., 2022) 0.386 $\pm$ 0.047 0.817 $\pm$ 0.011 0.329 $\pm$ 0.077 0.359 $\pm$ 0.054
HREP (Zhou et al., 2023) 0.578 $\pm$ 0.041 0.664 $\pm$ 0.017 0.468 $\pm$ 0.022 0.447 $\pm$ 0.061
ReCP (Li et al., 2024) 0.534 $\pm$ 0.057 0.804 $\pm$ 0.045 0.284 $\pm$ 0.076 0.325 $\pm$ 0.044
HAFusion (Sun et al., 2024) 0.631 $\pm$ 0.036 0.870 $\pm$ 0.010 0.613 $\pm$ 0.067 0.544 $\pm$ 0.035
RegionDCL (Li et al., 2023) 0.179 $\pm$ 0.053 0.402 $\pm$ 0.042 0.445 $\pm$ 0.041 0.190 $\pm$ 0.032
UrbanCLIP (Yan et al., 2024) 0.416 $\pm$ 0.006 0.186 $\pm$ 0.024 0.491 $\pm$ 0.003 0.288 $\pm$ 0.006
CityFM (Balsebre et al., 2024) 0.205 $\pm$ 0.018 0.614 $\pm$ 0.025 0.391 $\pm$ 0.027 0.271 $\pm$ 0.004
FlexiReg 0.766 $\pm$ 0.022 0.891 $\pm$ 0.024 0.753 $\pm$ 0.026 0.698 $\pm$ 0.014
Improvement 21.4% 2.4% 22.8% 28.3%
San Francisco Crime Check-in Service Call Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 0.594 $\pm$ 0.013 0.562 $\pm$ 0.021 0.479 $\pm$ 0.017   0.093 $\pm$ 0.016
MGFN (Wu et al., 2022) 0.601 $\pm$ 0.017 0.708 $\pm$ 0.010 0.468 $\pm$ 0.021   0.033 $\pm$ 0.048
HREP (Zhou et al., 2023) 0.612 $\pm$ 0.014 0.629 $\pm$ 0.032 0.461 $\pm$ 0.029   0.127 $\pm$ 0.023
ReCP (Li et al., 2024) 0.585 $\pm$ 0.075 0.783 $\pm$ 0.029 0.301 $\pm$ 0.119   0.067 $\pm$ 0.044
HAFusion (Sun et al., 2024) 0.682 $\pm$ 0.013 0.813 $\pm$ 0.024 0.612 $\pm$ 0.018   0.159 $\pm$ 0.006
RegionDCL (Li et al., 2023) 0.413 $\pm$ 0.021 0.437 $\pm$ 0.024 0.256 $\pm$ 0.024   0.027 $\pm$ 0.025
UrbanCLIP (Yan et al., 2024) 0.283 $\pm$ 0.014 0.334 $\pm$ 0.002 0.292 $\pm$ 0.008 -0.395 $\pm$ 0.005
CityFM (Balsebre et al., 2024) 0.334 $\pm$ 0.008 0.298 $\pm$ 0.008 0.396 $\pm$ 0.008   0.023 $\pm$ 0.003
FlexiReg 0.732 $\pm$ 0.014 0.859 $\pm$ 0.011 0.641 $\pm$ 0.021   0.480 $\pm$ 0.034
Improvement 7.3% 5.7% 4.7% 202%
4.3. Cross-country Applicability (Q2)
Table 3. Prediction Accuracy over Cities in Different Countries
Singapore Lisbon
Check-in Population Check-in Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
CityFM 0.239 $\pm$ 0.039 0.412 $\pm$ 0.010 0.195 $\pm$ 0.004 0.627 $\pm$ 0.003
Improvement 29.3% 41.0% 40.6% 16.6%
Table 3 reports the results (full results are in Appendix L). Our model FlexiReg outperforms all competitors consistently, with an improvement of at least 16.6% in $R^{2}$ over the best baseline models. These results confirm FlexiReg’s applicability across countries. Among the baseline models, CityFM performs better on Singapore, while UrbanCLIP is more suitable for Lisbon, underscoring the different urban characteristics of the two cities.
4.4. Adaptability to Region Formations (Q3)
(a) Crime
(b) Check-in
(c) Service Call
(d) Population
Figure 5. Adaptability to regions of varying sizes (NYC).
4.5. Ablation Study (Q4)
We study the effectiveness of FlexiReg model components with the following variants: (1) FlexiReg-w/o-PE removes the prompt enhancer from the downstream task learning stage. (2) FlexiReg-w/o-TAlign replaces the text-region alignment module with a direct concatenation of text and region embeddings. (3) FlexiReg-w/o-SVAlign replaces the street view-region alignment module with the summation of all street view image embeddings in a region, followed by concatenating these with the region embeddings. (4) FlexiReg-w/o-EC removes the environment context-based contrastive learning and directly use ResNet to generate visual embeddings for each street view image. (5) FlexiReg-w/o-CNN replaces the CNN branch with a GNN branch to process satellite images. (6) FlexiReg-w/o-Grid directly learns region embeddings without using the grid cells. (7) FlexiReg-w/o-FS uses all features (including textual descriptions and street view images) during cell embedding learning. (8) FlexiReg-w/o-WS replaces the weighted summation of cell embeddings with a direct summation.
Figure 6. Ablation study results (NYC).
4.6. Impact of Input Features (Q4)
Figure 7. Impact of input features (NYC).
4.7. Additional Results (Q5)
We conduct additional experiments with results reported in Appendix M, including model efficiency in Appendix M.1, model applicability to areas of different urban environments in Appendix M.2, impact of shape and size of the grid cells in Appendix M.3, and model sensitivity to key hyper-parameters in Appendix M.4.
5. Conclusion
Appendix A Related Work
These existing studies focus primarily on the general region embedding learning stage, while they neglect the downstream task learning stage. Prompting learning provides a promising solution for downstream task-based learning, which can guide the general region embeddings to adapt to specific tasks. Prompting learning has already achieved significant success in fields such as natural language processing (Schick and Schütze, 2021; Ben-David et al., 2022; Han et al., 2022) and computer vision (Zhou et al., 2022a, b; Khattak et al., 2023). However, in the context of region embedding learning, only one attempt has been made: HREP (Zhou et al., 2023), which generates random prompt embeddings and simply concatenates them with region embeddings for downstream tasks.
Discussion. Existing high-performing models rely heavily on human mobility data, while accessible features and their relevance to downstream tasks have been underexplored. We propose a novel, multimodal grid cell embedding learning module and an environment context-based contrastive learning approach to capture distinctive environmental characteristics and spatial correlations between different types of features. Besides, we introduce a prompt enhancer module to extract and integrate task-specific information into generic region embeddings to tailor for diverse downstream tasks.
Appendix B Notation
We summarize the frequently used symbols in Table 4.
@@ Appendix B Notation
Appendix C Input Cell Features
We detail our input features considered for each cell with the help of Fig. 8.
@@ Appendix C Input Cell Features
POI features. We use 15 representative POI categories to enhance feature representation, including educational institutions, commercial and industrial properties, accommodation, cultural and recreational venues, healthcare and medical facilities, entertainment venues, places of worship, food and drink establishments, parking facilities, transportation and transit facilities, residential properties, camping and outdoor recreation sites, sports and recreation facilities, financial services, and others.
Land use features. We use the pre-defined land use categories from OpenStreetMap, including grass, park, cemetery, forest, scrub, meadow, farmland, industrial, heath, retail, military, nature reserve, residential, commercial, orchard, farmyard, allotments, recreation ground, vineyard, and quarry.
Street view imagery features. To collect the street view images, we begin by sampling geocoordinate points along the road network at 100-meter intervals. Next, we remove redundant points that are within 20-meter to one another. In cases where the number of sampled points is less than 5 within a grid cell, we randomly sample additional points for that cell. Finally, we collect street view images from four different directions (0°, 90°, 180°, 270°) at each sampled point.
Appendix D Details of Grid-based Inter-view Feature Learning
We introduce a grid-based inter-view feature learning module to learn spatial correlations between grid cells across views. To learn the cross interactions between different views of the same cell, we simply apply a one-layer self-attention network (Vaswani et al., 2017) over $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$ which stacks intra-view cell embeddings $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$, and $\mathbf{Z}^{si}$. Now we denote $\mathbf{Z}$ and $\mathbf{Z}^{\mathcal{X}}$ as $\mathbf{Z}_{intra}$ and $\mathbf{Z}^{\mathcal{X}}_{intra}$, respectively, to distinguish with the embedding matrices used later for inter-view features, where $\mathcal{X}\in\{p,l,gn,si\}$.

### v3

# FlexiReg: Flexible Urban Region Representation Learning

@@ Abstract.
conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaccs: Information systems Location based servicesccs: Information systems Data mining
@@ 1. Introduction
Urban region representation learning has become increasingly popular in the community of urban computing (Zhang et al., 2017; Zheng et al., 2014; Yuan et al., 2024a; Su et al., 2024a, b; Xu et al., 2025; Wei et al., 2025; Chang et al., 2023, 2025), which aims to transform urban regions into vector representations, known as embeddings. These embeddings entail valuable insights on urban structures and properties, facilitating effective urban planning and management, such as designating functionalities for new development areas. They are also useful in various tasks related to daily life, such as crime count prediction (Wu et al., 2022; Sun et al., 2024; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024).
@@
Limitation 1. Existing methods heavily depend on mobility data and underutilize publicly accessible data. Human mobility data plays a critical role in learning effective region embeddings (Sun et al., 2024; Wu et al., 2022; Li et al., 2024; Zhang et al., 2020, 2023b). However, their limited availability, particularly in underdeveloped regions, together with privacy issues, prevent models using such data from a wider adoption. Recently, several studies (Li et al., 2023; Xi et al., 2022; Yan et al., 2024; Yong and Zhou, 2024) leverage features from publicly accessible data (e.g., POIs from OpenStreetMap) to enhance model applicability. However, these studies suffer from the effectiveness of the learning models. Their learned embeddings have reported lower accuracy for downstream tasks comparing with those learned by the mobility-based models.
@@
We summarize existing works for the issues above in Table 1.
@@
(2) It has a prompt learning process for its third stage, which enables it to flexibly utilize different types of features for different downstream tasks. We propose a novel prompt enhancer (PromptEnhancer) module to tailor region embeddings for downstream tasks by integrating textual and street-view imagery features. To capture task-relevant information, PromptEnhancer consists of a text-region alignment (T-RAlign) module and a street view-region alignment (SV-RAlign) module. T-RAlign incorporates task-specific geographic knowledge into region embeddings using dimension-wise similarity, while SV-RAlign extracts task-relevant visual features through adapted attention mechanisms. (Addressing Limitation 2)
@@ 2. Solution Overview
Model overview. Fig. 2 shows the overall structure of our model FlexiReg, which consists of three main learning stages. (1) FlexiReg takes a set of grid cells as the input. The grid cells come from a fine-grained partitioning over the spatial area of interest that we perform as part of data preparation (Section 3.1). FlexiReg learns the embeddings of grid cells across different features through a Multimodal Grid Cell Embedding Learning module (GridLearner, Section 3.2). (2) Then, the Adaptive Region Embedding Generation module (AdaRegionGen) aggregates the fine-grained cell embeddings to generate region embeddings for the input regions (Section 3.3). Starting from the fine-grained cell embeddings allows FlexiReg to flexibly adapt to different sets of regions which may come from different downstream tasks (or analytical tasks to explore different way to form regions). (3) Finally, the Prompt Enhancer module (PromptEnhancer) refines the generic region embeddings with extra features, guided by given tasks (Section 3.4).
@@ 3. Proposed Model
This section details the FlexiReg model. We summarize the frequently used symbols in Table 2.
@@
3.1. Data Preparation
3.1.1. Grid Cell Construction
3.1.2. Feature Preparation
We use six types of features for each cell, which are all publicly accessible, as shown in the Fig. 3
@@ 3.1.2. Feature Preparation
POI features. For each cell, we count the number of POIs that belong to one of 15 POI categories (educational institutions, commercial and industrial properties, accommodation, cultural and recreational venues, healthcare and medical facilities, entertainment venues, places of worship, food and drink establishments, parking facilities, transportation and transit facilities, residential properties, camping and outdoor recreation sites, sports and recreation facilities, financial services, and others.) from OpenStreetMap (OpenStreetMap, 2024) as the POI feature. We denote the POI feature of cell $c_{i}$ as $\mathbf{p}_{i}\in\mathbb{R}^{15}$.
Land use features. Similar to POI features, we count the numbers of zones that belong to 20 different land use types (grass, park, cemetery, forest, scrub, meadow, farmland, industrial, heath, retail, military, nature reserve, residential, commercial, orchard, farmyard, allotments, recreation ground, vineyard, and quarry.) within a cell. The land use feature of cell $c_{i}$ is denoted as $\mathbf{l}_{i}\in\mathbb{R}^{20}$.
Street view imagery features. Street view images capture finer-grained urban patterns. We use $\mathbf{sv}_{i}=\{\mathbf{sv}_{i,1},\mathbf{sv}_{i,2},\cdots\}$ to denote the set of street view images captured within the area of $c_{i}$, where each image is also in the shape of $\mathbb{R}^{H\times W\times 3}$. Note that different cells may have different numbers of street view images. To collect the street view images, we begin by sampling geocoordinate points along the road network at 100-meter intervals. Next, we remove redundant points that are within 20-meter to one another. In cases where the number of sampled points is less than 5 within a grid cell, we randomly sample additional points for that cell. Finally, we collect street view images from four different directions (0°, 90°, 180°, 270°) at each sampled point.
3.2. Multimodal Grid Cell Embedding Learning
3.2.1. Grid-based Intra-view Feature Learning
(1) $\alpha_{ij}^{g}$ $=\mathrm{Softmax}\left(\sigma\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{g}_{i}||\mathbf{W}\mathbf{z}^{g}_{j}||\mathbf{w}\mathbf{A}_{i,j}\right)\right)\right),$
(3) $\mathbf{z}^{si}_{i}=\mathrm{MLP}(\mathrm{ResNet}(\mathbf{si}_{i})),$
3.2.2. Grid-based Inter-view Feature Learning
Next, we introduce a grid-based inter-view feature learning module to learn spatial correlations between grid cells across views. To learn the cross interactions between different views of the same cell, we simply apply a one-layer self-attention network (Vaswani et al., 2017) over $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$ which stacks intra-view cell embeddings $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$, and $\mathbf{Z}^{si}$. Now we denote $\mathbf{Z}$ and $\mathbf{Z}^{\mathcal{X}}$ as $\mathbf{Z}_{intra}$ and $\mathbf{Z}^{\mathcal{X}}_{intra}$, respectively, to distinguish with the embedding matrices used later for inter-view features, where $\mathcal{X}\in\{p,l,gn,si\}$.